What is the pH of a buffer solution containing 0.63 NaHCO3 and 0.32 H2CO3


1
Expert's answer
2021-10-05T23:59:04-0400

pH=6.37 + Log"\frac{0.63}{0.32}\\"

= 6.37 + Log(1.96875)

=6.37 + 0.29419

=6.68419